# HAF-RM: A Hybrid Alignment Framework for Reward Model Training

### v2

@@ 5.2.1 Best-of-N
Figure 3: Win rate of responses selected by the HAF model compared to the baseline model.
We demonstrate the reliability of our trained reward model through Best-of-N pick, in which the reward model should pick the best one (the response with the highest reward) from several responses sampled from the same language model. The backbone for the reward model and the sampling model are the same, 8 and 4 responses are provided to the Mistral-based reward model the Phi-2-based reward model respectively, because Phi-2 is more likely to generate the same responses. The prompts for comparisons and ranking are listed in Appendix E, which reference AlpacaEval (Li et al., 2023b).
We report two evaluation metrics. Win rate: We use GPT-4-turbo to directly compare the responses from HaF reward model and baseline and report the win rate (Jang et al., 2023). Consistency with GPT: we use GPT-3.5-turbo to rank the sampled responses and calculate the recall of the top-1 and top-2 responses.
Top-1(%) Top-2(%)
HaF Baseline HaF Baseline
Phi-2 33.77 26.68 58.30 49.47
[…]
Mistral 13.31 11.55 25.27 23.49
[…]
Table 4: Top-k recall for HaF and the baseline. There are 4 candidate responses for Phi-2 and 8 for Mistral. The results are averaged over the recall values from all five datasets. The subscript “No harm” indicates that the result in that row is averaged over the AHP, CA, and harmless datasets instead of all datasets.
As shown in Figure 3 and Table 4, HaF demonstrates significant advantages over the baseline reward model in selecting responses especially for Phi-2 model in terms of both evaluation metrics. It is important to note that the average performance of the baseline reward model is comparable to random selection, suggesting that it has poor sensitivity and cannot effectively distinguish between responses when the quality differences are minimal. In contrast, the reward model obtained using HaF demonstrates good discriminative ability. Considering that the model can only learn to distinguish harmful from non-harmful responses from the BS and Harmless datasets, and that the responses generated by Phi-2 and Mistral are mostly harmless, we also report the average results on the remaining three datasets. When the safety-related datasets are excluded, both HAF and baseline show an improvement in average performance. Due to space limitations, the detailed results are presented in the appendix in Table 10.
@@ 5.2.2 RLHF
We also test HaF in the regular RLHF process: we train two reward models with HaF and the baseline method and then use them to train the policy models with RLHF. After training, GPT-3.5-turbo is introduced to compare the generations from the two policy models.
We conduct experiments using the Mistral model along with the AHP, CA, and Helpful datasets to investigate the reward model’s capability in optimizing for comprehensive preferences. Phi-2 is not used here as it shows great unstability during training which may not exhibit any performance improvement. Setups for reward model training and PPO are listed in Appendix A.
#Win #Lose Win rates(%)
AHP 285 215 57.00
CA 346 154 69.20
Helpful 243 256 48.70
Table 5: Win rates for the policy model trained with HaF reward model by RLHF.
HaF demonstrates a significant advantage on the AHP and CA datasets, while showing slightly worse performance compared to the baseline on the Helpful dataset. This indicates that the HaF reward model provides more effective guidance for the policy model. Given the widespread application of RLHF-like methods, HaF shows promising potential for active use in language model alignment in the near future. However, due to the simple experimental setup and the inherent instability of RLHF at small scales, the effectiveness of the HAF method in language model alignment still requires extensive exploration.
@@ 6 Related Work
Therefore, how to construct a model offering explicit preference feedback has naturally become a focal point of much research. To train a precise and robust reward model, many studies start from training with human preference data, and many works in the data field are largely centered around this. (Touvron et al., 2023) and (Zhao et al., 2022) provided different methods for using ranking data; (Wang et al., 2024) explored ways of measuring the strength of the data; while concerning datasets themselves, (Azar et al., 2023), (Knox et al., 2022) and (Hong et al., 2022) analyzed the impact of data preference strength on training from theoretical or practical perspectives. In addition, similar to the RAG technique (Lewis et al., 2020) in large language models, many methods (Li et al., 2023a; Sun et al., 2023) using external tools or references have also emerged, injecting new vitality into the development of reward models.
@@ 7 Conclusion
In this paper, we extend and improve the training framework of the current reward model. We split the training mechanism of the reward model into two stages: aligning model preference and optimizing the reward layer. Through introducing an additional constraint of policy loss, our hybrid alignment framework supervises the internal preference model at the token level while simultaneously optimizing the mapping layer at the seqneuce level, significantly improving the training effectiveness. We theoretically verify the validity of our method and demonstrate its reliability through systematic experiments.
@@ Appendix A Experiments Setup
lora rank 64 optimizer AdamW precision bf16
@@
To train the reward model, we use DPO Loss as the policy loss in HAF and set policy ratio $\alpha=0.2$. The learning rate is $1.0\times 10^{-5}$ for Phi-2 and Mistral-lora-baseline, $3.0\times 10^{-5}$ for Mistral-lora-HAF. A single RTX A6000 with 48GB memory is used for training the reward model. The model used for testing is the checkpoint that achieves the highest reward on the validation set.
For PPO training in Section 5.2.2, we utilize two RTX A6000 GPUs for parallel training with a total batch size of 4. The maximum number of new tokens generated is set to 128, and the learning rate is 1e-6. The training is conducted over a maximum of 20,000 episodes. We employ score scaling and score normalization and clip the scores between -3 and 3. All other settings follow the implementation in the TRL library. The model used for testing is the checkpoint that achieves the highest reward on the validation set. The generation config includes $top\_p=0.8$, $temperature=0.5$, $length\_penalty=1.3$, $repetition\_penalty=1.2$, $do\_sample=True$
@@ C.1 Deriving the Reward Loss Functions
In practice, there is no access to the ground truth reward of a response, so it is not applicable to solve the reward regression problem by directly optimizing the discrepancy between every predicted reward and the true reward. The Bradley-Terry model is introduced here to construct a solvable classification problem with one additional assumption – if one response is better than the other, then it wins with the probability of 100%. For a query $x$, a preferred response $y$ and a dispreferred response $y^{\prime}$, the predicted winning probability is $\mathrm{P}(y\succ y^{\prime})=\sigma(\bm{r}(x,y)-\bm{r}(x,y^{\prime}))$, and the ground truth $\mathrm{P}^{*}(y\succ y^{\prime})=\sigma(\bm{r}^{*}(x,y)-\bm{r}^{*}(x,y^{\prime}))=1$, so the standard reward loss is essentially a cross-entropy loss of the predicted winning probability.
$\mathcal{L}=$ $-\mathrm{P}^{*}(y\succ y^{\prime})\log\mathrm{P}(y\succ y^{\prime})$
$-\mathrm{P}^{*}(y^{\prime}\succ y)\log\mathrm{P}(y^{\prime}\succ y)$
$=$ $-\log\sigma(\bm{r}(x,y)-\bm{r}(x,y^{\prime}))$
The optimal model $\mathrm{F}^{*}$ and $\mathrm{\phi}^{*}$ are secretly hidden in the coefficient “1”.
@@ C.2 DPO as the Policy Loss
The derivation for policy loss is the same as reward loss in their essence. The policy model can be treated as a reward model with sequence probabilities reflecting the rewards (Rafailov et al., 2023, 2024). $reward(x,y)=\pi(x,y)/\pi_{ref}(x,y)$. With the Bradley-Terry model and the assumption of $\mathrm{P}(y\succ y^{\prime})=1$, DPO is also a legal loss function.
From this perspective, the DPO loss and reward loss share the same assumption of $P(y\succ y^{\prime})=1$. The reward model and the DPO-trained policy model are essentially doing the same task despite some formal differences (Rafailov et al., 2023, 2024). This may provide insight into why DPO is the most suitable among all policy losses.
Appendix D Mathematical Derivations
D.1 Inequality Scaling
@@
Here we get Prop. 1.
D.2 Derive the Final Inequality with the 3 Properties
@@
Based on these three properties, we can derive the result from Prop. 1.
Prop. 1
@@
which is Proposition 2.
Appendix E GPT Judgement
@@
Table 7 shows the consumption approximation for getting top-1, top-2 responses and the complete order out of 4/8 responses. We consider that performing a single sorting operation on eight responses with the model may result in a loss of precision. Besides, while binary comparisons exhibit high accuracy, repeated binary comparisons inevitably lead to cumulative errors and erroneous outcomes. Therefore, whether from a cost or accuracy standpoint, it is not a favorable option. In practice, we obtain the top 2 responses by ranking 4 responses with GPT-3.5-turbo at once. For 8 candidate responses, we first evenly divide them into two groups and use GPT to rank the responses of each group, then we rank the two sets of the top 2 responses to get the top 2 responses among 8 candidates.
@@
Phi-2HAF 28.68 52.51 32.69 53.35 37.52 66.21 45.44 74.26 24.52 45.15
Phi-2baseline 15.46 34.64 29.28 49.72 27.83 51.68 43.62 73.92 17.22 37.29
MistralHAF 17.42 31.22 9.94 17.70 16.00 28.81 13.68 27.57 9.50 21.07
Mistralbaseline 10.97 23.87 7.45 17.08 17.99 32.78 12.68 26.36 8.68 17.36

### v3

@@ 4.2.1 Best-of-N
Figure 4: Average win rates of responses selected by the HAF reward model, baseline model and the DPO reward model. Circles may overlap as different models select the same response.
We demonstrate the reliability of our trained reward model through Best-of-N selection, where the reward model should pick the best response (the one with the highest reward) from several responses sampled from the same generative model. The backbone for the reward model and the generation model is the same, with 8 and 4 responses are provided to the Mistral-Instruct reward model and the Phi-2 reward model, respectively. Because Phi-2 tends to generate more similar responses, reducing the need for 8 candidates. The prompts used for comparisons and ranking are listed in Appendix F, referencing AlpacaEval (Li et al., 2023b). We report two evaluation metrics. Win rate: We use GPT-4-turbo to rank the responses from HaF , DPO, and baseline reward model and report the win rate (Jang et al., 2023). Consistency: we use GPT-4-turbo to rank the sampled responses and calculate the recall of the top-1 and top-2 responses.
Phi-2 Mistral
All Chat All Chat
Top-1 Random 25.00 25.00 12.50 12.50
Baseline 27.43 28.97 16.03 18.27
DPO 22.94 26.39 12.81 13.85
HaF 33.77 37.19 18.19 21.12
Top-2 Random 50.00 50.00 25.00 25.00
Baseline 49.71 53.39 30.64 35.13
DPO 46.22 51.59 29.05 31.56
HaF 58.28 64.23 34.89 39.96
Table 4: Top-k recalls of different reward models. Random shows the recall when choosing responses randomly. The results are averaged over the recall values from all datasets. “Chat” indicates that the result in that column is averaged over the AHP, CA, and Harmless instead of all five datasets.
As shown in Figure 4 and Table 4, HaF demonstrates significant advantages over the baseline and DPO reward models in selecting responses in terms of both evaluation metrics, especially taking Phi-2 as the backbone. Notably, the recall scores of both DPO and baseline are close to those of random selection, indicating poor sensitivity and an inability to discern between responses with minimal quality differences. In contrast, the reward model trained by HaF exhibits good discriminative ability.
Considering that the model primarily learn to distinguish between harmful and non-harmful responses from the BS and Harmless datasets, and the responses generated by Phi-2 and Mistral are mostly benign, we also report average results on the remaining three datasets. When the safety-related datasets are excluded, all models show an improvement in average performance. The detailed results as well as the ArmoRM-judged results can be found in the appendix in Table 11, Figure 9.
Figure 4 presents the win rates of each method. We can observe that HAF consistently has the highest probability of selecting the best response (among the three methods), while DPO performs the worst. The frequency with which the baseline reward model and the HaF reward model select the same optimal response is considerably higher than their agreement with DPO. This difference is partly due to their modeling approaches: both HaF and the baseline reward model directly produce numerical rewards, whereas DPO derives rewards from token probabilities.
@@ 4.2.2 RLHF
We also test HaF in the standard RLHF process: we train two reward models respectively with HaF and the baseline method and then use them to train policy models through RLHF. After training, GPT-4 acts as the evaluator to compare the generations from the two policy models. We conduct two sets of experiments: one for training a Safety reward model using the BS and Harmless datasets; and the other for training a Chat reward model using the AHP, CA, and Helpful datasets. We compare the response quality of the policy models optimized after the same number of PPO steps by the baseline reward model and the HaF reward model.
Figure 5: Win rates for the policy model trained with the HaF reward model using RLHF compared to the baseline reward model, with each comparison made at the same training steps.
As shown in Table 5, the improvement of HaF is particularly evident on the Chat dataset, with its win rate increasing throughout the training, highlighting the superiority of the HaF reward model. In contrast, during safety training, the HaF reward model shows a significant advantage over the baseline model primarily in the middle stages of training. This is likely because both models have largely achieved harmless responses on the test set, resulting in minimal differentiation between the two reward models.
@@ 5 Related Work
Therefore, how to construct a model offering explicit preference feedback has naturally become a focal point of much research. To train a precise and robust reward model, many studies start from training with human preference data, and many works in the data field are largely centered around this. Touvron et al. (2023) and Zhao et al. (2022) provided different methods for using ranking data; Wang et al. (2024a) explored ways of measuring the strength of the data; while concerning datasets themselves, Azar et al. (2023), Knox et al. (2022) and Hong et al. (2022) analyzed the impact of data preference strength on training from theoretical or practical perspectives. In addition, similar to the RAG technique (Lewis et al., 2020) in large language models, many methods (Li et al., 2023a; Sun et al., 2023) using external tools or references have also emerged, injecting new vitality into the development of reward models.
@@ 6 Conclusion
In this paper, we extend and improve the training framework of the current reward model. We split the training mechanism of the reward model into two stages: aligning model preference and optimizing the reward layer. Through introducing an additional constraint of policy loss, our hybrid alignment framework supervises the internal preference model at the token level while simultaneously optimizing the mapping layer at the sequence level, significantly improving the training effectiveness. We theoretically verify the validity of our method and demonstrate its reliability through systematic experiments.
@@ Appendix A Experiments Setup
lora rank 64 optimizer AdamW precision fp16
@@
To train the reward model, we use DPO Loss as the policy loss in HAF and set policy ratio $\alpha=0.2$. The learning rate is $1.0\times 10^{-5}$ for Phi-2 and Mistral-Instruct with the baseline method, $3.0\times 10^{-5}$ for Mistral-Base and Mistral-Instruct using other methods, and the batch size is set at 16. These configurations are the optimal combination of learning rates (among $1.0\times 10^{-4}$, $3.0\times 10^{-5}$, $1.0\times 10^{-5}$, $3.0\times 10^{-6}$) and batch sizes (among $4$, $16$, $128$). A single RTX A6000 with 48GB memory is used for training the reward model. The model used for testing is the checkpoint that achieves the highest reward on the validation set.
For PPO training in Section 4.2.2, we set the total batch size at 16. The maximum number of new tokens generated is set to 256, and the learning rate is $2.0\times 10^{-5}$. The training is conducted over a maximum of 100,000 episodes. All other settings follow the implementation in the LLaMA-Factory library. The generation config includes top_p=$0.9$, do_sample=True.
@@ C.1 Deriving the Reward Loss Functions
In the Bradley-Terry model’s assumption, Oracle reward model outputs rewards in connection with the win rates:
$\underset{p\sim J}{\mathbb{E}}\mathbb{I}(y>y^{\prime};x,p)=-\log\sigma[\bm{r}^{*}(x,y)-\bm{r}^{*}(x,y^{\prime})]$ (6)
where $p$ is a judge (annotator) sampled from the judge distribution $J$.
As we only focus on the reward differences between responses to the same prompt, there exists another metric denoted as $\mathrm{D}_{1}^{\prime}$ for calculating the reward loss:
$\mathcal{L}_{s}=\underset{x,y,y^{\prime}}{\mathbb{E}}\mathrm{D}_{1}^{\prime}[$ $r(x,y)-r(x,y^{\prime}),$
$\quad r^{*}(x,y)-r^{*}(x,y^{\prime})]$
As $-\log\sigma(\cdot)$ is monotonically increasing, so there exists a metric $\mathrm{D}_{1}^{\prime\prime}$, such that
$\mathrm{D}_{1}^{\prime}[r(x,y)-r(x,y^{\prime}),r^{*}(x,y)-r^{*}(x,y^{\prime})]$
$=$ $\mathrm{D}_{1}^{\prime\prime}[-\log\sigma(\bm{r}(x,y)-\bm{r}(x,y^{\prime})),$
$\qquad\qquad\qquad\ -\log\sigma(\bm{r}^{*}(x,y)-\bm{r}^{*}(x,y^{\prime}))]$
$=$ $\mathrm{D}_{1}^{\prime\prime}[-\log\sigma(\bm{r}(x,y)-\bm{r}(x,y^{\prime})),$
$\qquad\qquad\qquad\qquad\qquad\underset{p\sim J}{\mathbb{E}}\mathbb{I}(y>y^{\prime};x,p)]$
Let $\mathrm{D}_{1}^{\prime\prime}$ be the cross-entropy loss, and let $\mathrm{P}(x,y,y^{\prime})=-\log\sigma(\bm{r}(x,y)-\bm{r}(x,y^{\prime}))$,
$\mathcal{L}_{s}=\underset{x,y,y^{\prime}}{\mathbb{E}}[\mathrm{P}(x,y,y^{\prime})\cdot\underset{p\sim J}{\mathbb{E}}\mathbb{I}(y>y^{\prime};x,p)$
$\quad+(1-\mathrm{P}(x,y,y^{\prime}))\cdot(1-\underset{p\sim J}{\mathbb{E}}\mathbb{I}(y>y^{\prime};x,p))]$
$=\underset{p\sim J}{\underset{x,y,y^{\prime}}{\mathbb{E}}}[\mathrm{P}(x,y,y^{\prime})\cdot\mathbb{I}(y>y^{\prime};x,p)$
$\quad+(1-\mathrm{P}(x,y,y^{\prime}))\cdot(1-\mathbb{I}(y>y^{\prime};x,p))]$
which is exactly Eq. 2 when we sample from $\mathcal{D}$.
@@ C.2 DPO as the Policy Loss
The derivation for policy loss is the same as reward loss in their essence. The policy model can be treated as a reward model with sequence probabilities reflecting the rewards (Rafailov et al., 2023, 2024). $reward(x,y)=\log[\pi(x,y)/\pi_{ref}(x,y)]$.
From this perspective, the DPO loss and reward loss share the same assumption (Eq. 6). The reward model and the DPO-trained policy model are essentially doing the same task despite some formal differences (Rafailov et al., 2023, 2024).
Appendix D Mathematical Enlightenment
D.1 Theoretical Explanation for the Claims
Inequality for claim 1.
Unless $K$ can exactly fit $K^{*}$, there exists $\epsilon>0$, such that
holds for all $\alpha\in(0.1,2)$, where $K_{H},\phi_{H}=\underset{\mathrm{K},\phi}{\mathrm{argmin}}\mathcal{L}_{H}$ in Equation 5 and $\phi_{s}=\underset{\phi}{\mathrm{argmin}}\mathcal{L}_{s}$ in Equation 2. Here we use $\mathrm{argmin}$ to represent the best models optimized with the corresponding loss functions, so $\phi_{H}$ and $\phi_{s}$ are not equal to $\phi^{*}$ although $\phi^{*}$ is the minimum mathematically.
Inequality for claim 2.
We obtain informally here an upper bound on the model preference error. By tuning the hyperparameter $\alpha$, the right term can be strictly negative.
D.2 Inequality Scaling
@@
Here we get the first inequality.
D.3 Derive the Final Inequality with the 3 Properties
@@
Based on these three properties, we can derive the result from Appendix D.1.
Inequality 2
@@
Appendix E Experiment Results
E.1 Consistency
Figure 7: The distribution of the reward model’s and DPO model’s outputs on test data when trained with identical data.
The x-axis in Figure 2 represents the reward difference between the responses generated by the DPO model and those generated by the HAF model’s policy head. This difference is scored by the reward model trained on the same data distribution, which we refer to as the Oracle reward model. We retain the checkpoints from the training processes of both DPO and HaF model and identify potential model pairs with similar performance using five methods (corresponding to the five colors in the figure). This similarity in performance ensures that the higher reward is not a result of better response quality. The five methods include “reward” (similar scores from the Oracle reward model), “acc” (similar binary classification accuracy), “loss” (similar loss values), “margin” (similar average margins of model predictions), and “step” (same training steps).
It can be observed that the differences in HaF scores are almost always higher than those from the Oracle reward model. This suggests that the preferences of the reward model are influenced by the preferences of the shared parameter policy model, providing some evidence for the existence of an Internal Preference Model.
Also shown in Figure 7, we independently trained a DPO model and a reward model using the same data and observed a strong positive correlation (even linearity) in their predictions on the test data. This indicates a significant similarity in the preference modeling processes of the DPO model and the reward model. A response preferred by the reward model will also be preferred by the DPO model, which we introduce the concept of the “Internal Preference Model” to explain.
E.2 Overall Performance
Model Metric Helpful Harmless CA BS AHP
Phi-2 ppwin 0.74 1.00 0.60 0.74 2.52
pplose 0.92 0.97 1.09 0.60 2.55
Mistral-base ppwin 0.42 0.65 0.51 0.38 0.75
pplose 0.62 0.63 0.87 0.28 0.98
Mistral-Instruct ppwin 3.50 5.13 2.33 1.58 1.98
pplose 6.08 5.81 3.52 1.31 2.67
Table 6: Variances of the corresponding metrics. “pp” means token-wise perplexity. The subscript “win” refers to the better response while “lose” refers to the worse response.
ppwin pplose ppwin-pplose
$corr$ $-0.8166$ $-0.9492$ $-0.9064$
$p$ $0.0916$ $0.0136$ $0.0339$
Table 7: The Pearson correlation coefficient between the variance of the token-wise perplexity of Mistral-Instruct and the difference in accuracy between the reward model trained with DPO and the accuracy of the baseline training. “corr” indicates the Pearson correlation coefficient, while “p” indicates significance.
Table 6 shows the token-wise perplexity calculated by each model for each dataset.
$pp=-\frac{\log\mathrm{Prob}(sequence)}{\mathrm{Length}(sequence)}$
Another interesting finding is that the variance of the token-wise perplexity (pp) values for Mistral-Instruct shows a very strong negative correlation with the performance of the DPO reward model. Table 7 calculates the Pearson correlation coefficient between the variance of the pp values and the performance difference between the DPO reward model and the baseline reward model, indicating that this negative correlation is highly significant. This finding may provide valuable insights for aligning well-trained (but not yet well-aligned) models.
E.3 Best of N
In Table 11 we list the recall value on each dataset. We show in Figure 8 and Figure 9 the win rates on each dataset judged by gpt-4-turbo-2024-04-09 and ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024b), respectively.
(a) AHP
(b) BS
(c) CA
(d) Helpful
(e) Harmless
Figure 8: Win rates on each dataset judged by gpt-4-turbo-2024-04-09
(a) AHP
(b) BS
(c) CA
(d) Helpful
(e) Harmless
(f) Avg.
Figure 9: Win rates on each dataset judged by ArmoRM-Llama3-8B-v0.1
Appendix F GPT Judgement
@@
Table 8 shows the consumption approximation for getting top-1, top-2 responses and the complete order out of 4/8 responses. We consider that performing a single sorting operation on eight responses with the model may result in a loss of precision. Besides, while binary comparisons exhibit high accuracy, repeated binary comparisons inevitably lead to cumulative errors and erroneous outcomes. Therefore, whether from a cost or accuracy standpoint, it is not a favorable option. In practice, we obtain the top 2 responses by ranking 4 responses with gpt-4-turbo-2024-04-09 at once. For 8 candidate responses, we first evenly divide them into two groups and use GPT to rank the responses of each group, then we rank the two sets of the top 2 responses to get the top 2 responses among 8 candidates.
@@
Phi-2HAF 28.67 52.51 32.49 53.06 37.46 65.94 45.44 74.25 24.79 45.67
Phi-2DPO 20.85 45.62 17.68 39.71 31.89 56.07 26.42 53.07 17.87 36.67
Phi-2baseline 15.45 34.63 32.85 50.90 27.84 51.64 43.62 73.91 17.41 37.48
MistralHAF 22.86 41.95 14.61 25.38 24.12 42.69 16.39 35.24 12.99 29.23
MistralDPO 14.57 32.91 10.00 24.61 13.68 30.62 13.31 31.14 12.52 25.98
Mistralbaseline 14.82 30.40 13.07 24.61 25.05 43.85 14.95 31.14 12.29 23.20